P(A)= 0.3
P(B/A)= 0.6
P(A and B)= ?

Answer by stanbon(75887)

P(A)= 0.3
P(B/A)= 0.6
P(A and B)= P(B|A)*P(A)
= 0.6*0.3 = 0.18
Cheers,
Stan H.